# Optimal quantum communication networks: capacitance versus security

### v1

@@ II.3.2 Structure of MQE networks
Figure 4: (a) Density $\rho$ as a function of $\alpha$, for $p=1-1/e$ and two values of $L/\lambda_{0}$ Averages over $100$ realizations for $N=128~{}(\text{squares}),N=256~{}(\text{circles})$, and over $10$ realizations for $N=512~{}(\text{triangles}),N=1024~{}(\text{diamonds})$. Dashed lines correspond to the values of $\alpha_{c}^{0\to 1}$ (red), $\alpha_{c}^{1\to 2}$ (green) and $\alpha_{c}^{2\to 3}$ (purple) evaluated using Eq. (12). (b) Dependence of the density on $N$ for $L/\lambda_{0}=0.1$. (c) Dependence of the density on $N$ for $L/\lambda_{0}=10$.
@@
Figure 5: Plot of $\mathcal{L}^{*}$ vs $N$ for several values of $\alpha$, $p=1-1/e$, (a) $L/\lambda_{0}=0.1$, (b) $L/\lambda_{0}=10$.
The optimal network can be also characterized by measuring a quantity analogous to the betweenness usually considered in network analysis. The betweenness [44] of a node is the number of shortest paths among all pairs of nodes in the network that go through that node. For our purposes it is useful to define a modified betweenness where, instead of topological shortest paths, optimal paths are considered. Such a quantity provides a measure of the relevance of users, i.e., how crucial is their presence (and how damaging their removal). In Fig. 6 we plot the histogram of the number of nodes having a given modified betweenness for various $\alpha$. For $\alpha>\alpha_{c}$ the optimal network is fully connected and the values of the betweenness are all zero. When $\alpha$ becomes smaller than $\alpha_{c}$ a homogeneous distribution appears. As $\alpha$ is progressively reduced the distribution gets more heterogeneous, becoming extremely broad for $\alpha\to 0$. In such a case some nodes are particularly crucial and the network is overall highly vulnerable to external attacks.
Figure 6: Distribution of the number of nodes having a given modified betweenness for $N=64$, $p=1-1/e$, $L/\lambda_{0}=0.1$ (a), $L/\lambda_{0}=10$ (b), and several values of $\alpha$.
@@ III.1 Optimization algorithm
The task of constructing MQE networks cannot be addressed using dynamic programming methods such as Dijkstra’s [41], Prim’s [45] or Pollack’s [42] algorithms. Specifically, because of the inherently nonlocal nature of the problem arising from the interplay of topological and geometric terms in the efficiency functional, if the optimal path between $a$ and $b$ goes through node $c$ nothing guarantees that the subpath between $a$ and $c$, belonging to $\left\{a\rightarrow b\right\}^{*}$, yields also the optimal path between $a$ and $c$. Consequently, we have to resort to alternative approaches to tackle this problem. In Ref. [42] Pollack presents, among others, a matrix-based algorithm for solving the maximum capacity route problem, which is equivalent for us to finding the maximum capacitance between any pair of users. Starting from a matrix whose element $Q^{(0)}_{jk}$ is the capacitance of the link between node $j$ and $k$ given by Eq. (1) and $Q^{(0)}_{jj}=\infty$, an iterative procedure is defined. The elements of $Q^{(m)}$ at step $m$ are defined by
@@
where $\ell=1,\ldots,N$. In this way the element $Q^{(m)}_{jk}$ is the maximum capacitance of paths between $j$ and $k$ going through at most $m$ intermediate nodes. Iterating, the process convergence is reached at most when $m+1=N-1$. Individual elements of $Q^{(m)}$ provide at convergence the maximum capacitance between all node pairs. We modify Pollack’s original algorithm as follows. Given $\alpha$, at each iteration $m$ we evaluate the quantity
$\epsilon_{jk}=(1-\alpha)Q^{(m)}_{jk}+\alpha\log(1-p)m$ (15)
This is the maximal efficiency of paths between $j$ and $k$ of length at most $m+1$. For each pair $(j,k)$ the maximum of this quantity as a function of $m$ defines a value $m^{*}_{jk}$. Note that $m^{*}$ needs not to be the same for all pairs. The efficiency of the optimal path between $j$ and $k$ is thus
@@
This procedure gives the efficiency of the optimal paths between any two users. In order to construct one of these (in principle many) paths we proceed as follows. The capacitance $Q^{(m^{*}_{jk})}_{jk}$ associated to the optimal path between $j$ and $k$ necessarily appears in the initial matrix $Q^{(0)}$, as element $Q^{(0)}_{ln}$. All entries of $Q$ are assumed to be different. This indicates that the optimal path between $j$ and $k$ necessarily goes through the link between $l$ and $n$, which, indeed, determines the capacitance of the optimal path. Starting from the matrix $Q^{(0)}$ we construct the unweighted graph $A$ containing an edge between all node pairs such that the corresponding element in $Q^{(0)}$ is larger than $Q^{(0)}_{ln}$. In this graph $A$ we find the optimal path as the topological shortest path going from $j$ to $k$ with the constraint that it goes through the link $(l,n)$. Clearly, although the link between $l$ and $n$ is uniquely determined there are often many alternative shortest paths going through it; this gives rise to a large number of degenerate optimal paths.
@@ III.2 The efficiency of the fully connected network
$Q_{FC}=-\frac{\langle\ln(z/\tilde{\lambda}\rangle)}{\ln{2}}.$ (20)
@@
$Q_{FC}\approx\frac{2\pi}{\ln{2}}\left(\frac{\lambda_{0}}{L}\right)^{2}.$ (23)
In this regime, the value of $Q_{FC}$ is determined by the contribution of the few links shorter than $\lambda_{0}$, which have a large capacitance.
@@ IV Conclusions
The realization of large-scale quantum communication networks is a task of crucial relevance for quantum cryptography and quantum computing applications. Existing quantum network implementations employ intermediate trusted nodes as a practical and efficient means to connect remote users, thereby overcoming the limitations imposed by rate/distance bounds. Relying on trusted nodes, however, carries the inherent risk associated with the probability of encountering malicious nodes. Given these constraints, in this work we addressed the problem of designing optimal quantum communication networks connecting a set of users randomly distributed in a square of size $L$. For each pair of users, we determined the optimal path connecting them by maximizing the quantum efficiency of the path and we constructed the optimal network, called maximal quantum efficiency network as the graph union of the optimal paths. MQE networks therefore provide the optimal balance between security and quantum communication rate and they interpolate between Maximum Spanning Trees and Fully Connected networks, representing, respectively, the topologies having maximum quantum capacitance and maximum security. We carefully analyzed the performance of the MQE networks showing that the optimization can largely increase the average capacitance while keeping high levels of security. We also analyzed structural properties of MQE networks by means of numerical and analytical methods, showing how the tradeoff between capacitance and security affects their topological properties.
We expect the performance of MQE networks to be significantly affected by the spatial distribution of points; for simplicity we assumed a uniform distribution but it would be interesting to consider other possibilities. Another crucial assumption underpinning our work is that of single-path routing; more powerful routing strategies, where systems are transmitted in parallel through different quantum communication channels have been proposed to improve the capacitance [13] or the security [40]. In these situations, an extension of our algorithm may provide a way to combine the two approaches to fulfill simultaneously well-defined capacitance and security requirements. The degeneracy of the optimal path could then acquire further practical relevance. In conclusion, we stress that our freely available code [48] can be straightforwardly used to design optimal quantum communication networks for given location of users
@@ Acknowledgements
This study was funded by European Union - PON Ricerca e Innovazione 2014-2020 FESR /FSC - Project ARS01_$\_00734$ QUANCOM, Ministero dell’Università e Ricerca - Project PNRR MUR PE_$\_0000023$-NQSTI and PNRR MUR project CN_$\_00000013$-ICSC. The funders played no role in study design, data collection, analysis and interpretation of data, or the writing of this manuscript.

### v3

@@ II.3.2 Structure of MQCE networks
Figure 4: The link density of the optimal networks. (a) Density $\rho$ as a function of $\alpha$, for $p=1-1/e$ and two values of $L/\lambda_{0}$ Averages over $100$ realizations for $N=128~{}(\text{squares}),N=256~{}(\text{circles})$, and over $10$ realizations for $N=512~{}(\text{triangles}),N=1024~{}(\text{diamonds})$. Dashed lines correspond to the values of $\alpha_{c}^{0\to 1}$ (red), $\alpha_{c}^{1\to 2}$ (green) and $\alpha_{c}^{2\to 3}$ (purple) evaluated using Eq. (12). (b) Dependence of the density on $N$ for $L/\lambda_{0}=0.1$. (c) Dependence of the density on $N$ for $L/\lambda_{0}=10$.
@@
Figure 5: Size dependence of the average length of optimal paths. Plot of $\mathcal{L}^{*}$ vs $N$ for several values of $\alpha$, $p=1-1/e$, (a) $L/\lambda_{0}=0.1$, (b) $L/\lambda_{0}=10$.
The optimal network can be also characterized by measuring a quantity analogous to the betweenness usually considered in network analysis. The betweenness [45] of a node is the number of shortest paths among all pairs of nodes in the network that go through that node. For our purposes it is useful to define a modified betweenness where, instead of topological shortest paths, optimal paths are considered. Such a quantity provides a measure of the relevance of users, i.e., how crucial is their presence (and how damaging their removal). Nodes with high betweenness have, just because of their position in the topology, a high impact on the security of the network. In Fig. 6 we plot the histogram of the number of nodes having a given modified betweenness for various $\alpha$. For $\alpha>\alpha_{c}$ the optimal network is fully connected and the values of the betweenness are all zero. When $\alpha$ becomes smaller than $\alpha_{c}$ a homogeneous distribution appears. As $\alpha$ is progressively reduced the distribution gets more heterogeneous, becoming extremely broad for $\alpha\to 0$. In such a case some nodes are particularly crucial and the network is overall highly vulnerable to external attacks.
Figure 6: The impact of individual nodes on security is heterogeneous. Distribution of the number of nodes having a given modified betweenness for $N=64$, $p=1-1/e$, $L/\lambda_{0}=0.1$ (a), $L/\lambda_{0}=10$ (b), and several values of $\alpha$. The lower $\alpha$, the broader the distribution becomes. Black dashed lines are guides to the eye.
@@ III.1 Optimization algorithm
The task of constructing MQCE networks cannot be addressed using dynamic programming methods such as Dijkstra’s [42], Prim’s [46] or Pollack’s [43] algorithms. Specifically, because of the inherently nonlocal nature of the problem arising from the interplay of topological and geometric terms in the communication efficiency functional, if the optimal path between $a$ and $b$ goes through node $c$ nothing guarantees that the subpath between $a$ and $c$, belonging to $\left\{a\rightarrow b\right\}^{*}$, yields also the optimal path between $a$ and $c$. Consequently, we have to resort to alternative approaches to tackle this problem. In Ref. [43] Pollack presents, among others, a matrix-based algorithm for solving the maximum capacity route problem, which is equivalent for us to finding the maximum capacitance between any pair of users. Starting from a matrix whose element $Q^{(0)}_{jk}$ is the capacitance of the link between node $j$ and $k$ given by Eq. (1) and $Q^{(0)}_{jj}=\infty$, an iterative procedure is defined. The elements of $Q^{(m)}$ at step $m$ are defined by
@@
where $\ell=1,\ldots,N$. In this way the element $Q^{(m)}_{jk}$ is the maximum capacitance of paths between $j$ and $k$ going through at most $m$ intermediate nodes. Iterating, the process, convergence is reached at most when $m+1=N-1$. Individual elements of $Q^{(m)}$ provide at convergence the maximum capacitance between all node pairs. We modify Pollack’s original algorithm as follows. Given $\alpha$, at each iteration $m$ we evaluate the quantity
$\epsilon_{jk}=(1-\alpha)Q^{(m)}_{jk}+\alpha m\log(1-p).$ (15)
This is the maximal communication efficiency of paths between $j$ and $k$ of length at most $m+1$. As $m$ is increased, the first term in Eq. (15) grows and tends to a constant, while the second is negative and is linear in $m$. Hence, for each pair $(j,k)$, $\epsilon_{jk}$ reaches a maximum as a function of $m$ for a value $m^{*}_{jk}$. Note that $m^{*}$ needs not to be the same for all pairs. The communication efficiency of the optimal path between $j$ and $k$ is thus
@@
This procedure gives the communication efficiency of the optimal paths between any two users. In order to construct one of these (in principle many) paths we proceed as follows. The capacitance $Q^{(m^{*}_{jk})}_{jk}$ associated to the optimal path between $j$ and $k$ necessarily appears in the initial matrix $Q^{(0)}$, as element $Q^{(0)}_{ln}$. All entries of $Q$ are assumed to be different. This indicates that the optimal path between $j$ and $k$ necessarily goes through the link between $l$ and $n$, which, indeed, determines the capacitance of the optimal path. Starting from the matrix $Q^{(0)}$ we construct the unweighted graph $A$ containing an edge between all node pairs such that the corresponding element in $Q^{(0)}$ is larger than $Q^{(0)}_{ln}$. In this graph $A$ we find the optimal path as the topological shortest path going from $j$ to $k$ with the constraint that it goes through the link $(l,n)$. Clearly, although the link between $l$ and $n$ is uniquely determined there are often many alternative shortest paths going through it; this gives rise to a large number of degenerate optimal paths.
@@ III.2 The efficiency of the fully connected network
$Q_{\text{FC}}=-\frac{\langle\ln(z/\tilde{\lambda}\rangle)}{\ln{2}}.$ (20)
@@
$Q_{\text{FC}}\approx\frac{2\pi}{\ln{2}}\left(\frac{\lambda_{0}}{L}\right)^{2}.$ (23)
In this regime, the value of $Q_{\text{FC}}$ is determined by the contribution of the few links shorter than $\lambda_{0}$, which have a large capacitance.
@@ IV Conclusions
The realization of large-scale quantum communication networks is a task of crucial relevance for quantum cryptography and quantum computing applications. Existing quantum network implementations employ intermediate trusted nodes as a practical and efficient means to connect remote users, thereby overcoming the limitations imposed by rate/distance bounds. Relying on trusted nodes, however, carries the inherent risk associated with the probability of encountering malicious nodes. Given these constraints, in this work we addressed the problem of designing optimal quantum communication networks connecting a set of users randomly distributed in a square of size $L$. For each pair of users, we determined the optimal path connecting them by maximizing the quantum communication efficiency of the path and we constructed the optimal network, called maximal quantum communication efficiency network as the graph union of the optimal paths. MQCE networks therefore provide the optimal balance between security and quantum communication rate and they interpolate between Maximum Spanning Trees and Fully Connected networks, representing, respectively, the topologies having maximum quantum capacitance and maximum security. We carefully analyzed the performance of the MQCE networks showing that the optimization can largely increase the average capacitance while keeping high levels of security. We also analyzed structural properties of MQCE networks by means of numerical and analytical methods, showing how the tradeoff between capacitance and security affects their topological properties. Our work proposes a systematic and scalable approach for quantum communication network optimization that relies on the construction of a network model and the corresponding quantum communication efficiency functional. So far we considered simple networks featuring only trusted nodes and lossy bosonic links but our work lays the basis for the study of general quantum communication networks.
We expect the performance of MQCE networks to be significantly affected by the spatial distribution of points; for simplicity we assumed a uniform distribution but it would be interesting to consider other possibilities. A further crucial assumption underpinning our work is that of single-path routing; more powerful routing strategies, where systems are transmitted in parallel through different quantum communication channels have been proposed to improve the capacitance [13, 49] or the security [41]. In these situations, an extension of our algorithm may provide a way to combine the two approaches to fullfill simultaneously well-defined capacitance and security requirements. The degeneracy of the optimal path could then acquire further practical relevance. Eventually, our method could be extended to different quantum key distribution schemes such as entanglement-based quantum key distribution [50, 51, 52]. In conclusion, we stress that our freely available code [53] can be straightforwardly used to design optimal quantum communication networks for given location of users.
@@ Acknowledgements
This work was funded by the HEISINGBERG “Spatial Quantum Optical Annealer for Spin Hamiltonians” EU Research and Innovation Project under Grant Agreement No. 101114978. V.B. acknowledge support from Project PNRR MUR PE_$\_0000023$-NQSTI financed by the European Union Next Generation EU. V.B. and L.P. acknowledge support from PON Ricerca e Innovazione 2014-2020 FESR /FSC - Project ARS01_$\_00734$ QUANCOM, Ministero dell’Università e Ricerca and PNRR MUR project CN_$\_00000013$-ICSC financed by the European Union Next Generation EU. The funders played no role in study design, data collection, analysis and interpretation of data, or the writing of this manuscript.